μ=220 grams σ=25 grams


If a mango is selected at random, what is the probability that its weight lies between 180 and 250 grams?

1
Expert's answer
2022-05-06T12:43:54-0400

We have a normal distribution, "\mu=220, \sigma=25."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"z_1=\cfrac{180-220}{25}=-1.60;\\
z_2=\cfrac{250-220}{25}=1.20;\\
P(180<X<250)=\\
=P(-1.60<Z<1.20)=\\
=P(Z<1.20)-P(Z<-1.60)=\\
=0.8849-0.0548=0.8301\\
\text{ (from z-table)}."